OpenStudy (anonymous):

What polynomial has roots of -1, -2, and 3 ? Will fan and give medal.

OpenStudy (anonymous):

A. x3 + 6x2 + 11x + 6 B. x3 - x2 - 5x + 2 C. x3 + 2x2 - 7x + 2 D. x3 - 7x - 6

OpenStudy (anonymous):

\[\alpha \beta \gamma=-\frac{D}{A}\] where \[\alpha, \beta, \gamma\] are your roots and \[A\] is the coefficient of the cubic term, \[D\] is the constant term. Multiply the roots simply and verify which polynomial is satisfying it

OpenStudy (anonymous):

That doesn't make sense to me.

OpenStudy (amorfide):

or you could just substitute in the values given, x=-1, x=-2, x=3 see which equation becomes zero when you substitute those values in

OpenStudy (anonymous):

That is a formula which relates roots and coefficients of a polynomial. For a quadratic polynomial you may have learned these formulae: \[\alpha \beta=\frac{C}{A}\]\[\alpha+\beta=-\frac{B}{A}\] For cubic you have \[\alpha \beta \gamma=-\frac{D}{A}\]\[\alpha+\beta+\gamma=-\frac{B}{A}\]\[\alpha \beta+\beta \gamma+\alpha \gamma=\frac{C}{A}\]

OpenStudy (anonymous):

You can also put the value of x, true but then you'd have to check all 3 roots

OpenStudy (anonymous):

I'm on a timed test, I'm just gonna guess.